OpenStudy (anonymous):

Simplify using properties of logarithms. log(x) - 1/2log(y)

mathslover (mathslover):

Is it : \(\log(x) - \cfrac{1}{2} \log (y)\) ?

OpenStudy (anonymous):

yes

mathslover (mathslover):

Fine! So, use the prop. we used in the previous question for the second term - \( a \log x = \log (x^a) \)

mathslover (mathslover):

So, what is : \(\cfrac{1}{2} \log (y)\) ?

OpenStudy (anonymous):

log(y^1/2)

mathslover (mathslover):

Right, so, we get : \(\log (x) - \log (y ^{\cfrac{1}{2}})\)

mathslover (mathslover):

Now, we will use : \(\log a - \log b = \log (\cfrac{a}{b})\)

mathslover (mathslover):

So, can you simplify the above expression further using the above property.

OpenStudy (anonymous):

log(x) / log(y^1/2)

mathslover (mathslover):

Again , no! You have to put a/b in a single log.

OpenStudy (anonymous):

Oh I always hit a bump on that

mathslover (mathslover):

You need to take care of that. It can harm you in future! Everything else is fine with you :) Just improve this mistake, and you will rock the exams ... :)

OpenStudy (anonymous):

log(x / y^1/2)

mathslover (mathslover):

That's right :) Good work. You can rationalize it further (the inner fraction) \(\log (\cfrac{x}{\sqrt{y}}) = \log (\cfrac{x}{\sqrt{y}} \times \cfrac{\sqrt{y}}{\sqrt{y}} ) = \log (\cfrac{x\sqrt{y}}{y})\) Though, the origial one is good too!

OpenStudy (anonymous):

What is the trick to remember all the properties of logarithms? I'm going to put them down on my one page of notes for the exam today

mathslover (mathslover):

Well, the best trick to remember formulas/properties in Mathematics is to practice a lot. If you practice a lot of questions , then you'll never forget them - that's my personal experience. For example, you just did a logarithm question like this, and if you do 2-3 more questions like this one, then you will remember the prop. we used here forever. You'll not forget them easily.

OpenStudy (anonymous):

We used the product rule and the quotient rule.

mathslover (mathslover):

Right... !

mathslover (mathslover):

Also, not to forget power rule.

OpenStudy (anonymous):

I got it!